# Universal Kerr-thermal dynamics of self-injection-locked microresonator dark pulses

### v1

@@
We further perform numerical simulation of Eq. (1–5) using experimentally obtained or realistic values for each parameter. Details are found in Supplementary Information Note 3. Figure 2(d,e) presents the simulated output optical power and frequency, which agree with the experimental data on both the overall trend (due to the photo-thermal effect) and the platicon steps (due to the Kerr nonlinearity). Figure 2f shows the zoom-in profile of the gray-shaded zoom in Fig. 2(d,e), and highlights the full detuning range of platicon steps with gray curves. Steps corresponding to those in Fig. 2c are marked. We emphasize that, previous efforts Kondratiev et al. (2020); Lihachev et al. (2022); Wang et al. (2022) to model platicon formation in the SIL regime fail to investigate and analyze the transient behaviour, and thus fail to reveal the step features. Meanwhile, they have also omitted the photo-thermal effect.
@@
Figure 3b presents the simulated platicon spectra for each labelled steps in Fig. 2f. They conform the experimental data in Fig. 3a, not only on the spectral envelopes but also on the number of fringes (marked with arrows). Figure 3c presents the simulated time-domain pulse shapes, revealing that the fringes are related to the oscillating tails at the bottom of the dark pulse (marked with arrows). Unlike the steps of bright dissipative solitons Guo et al. (2017); Voloshin et al. (2021), we demonstrate here that platicon steps can occur even in the single-platicon state, i.e. only one dark pulse in the microresonator.
@@
Low-noise photonic microwave generation. Finally, we showcase an immediate application of our findings, i.e. to generate low-noise photonic microwave which are ubiquitously deployed in modern information systems for wireless communication, timing and radar. Photonic microwave – microwave synthesized via photonics – allows unrivalled noise performance and bandwidth breaking the bottleneck of their electronic counterparts Fortier et al. (2011); Li et al. (2014); Xie et al. (2016). Among various photonics approaches, microcombs, which coherently channel microwave and terahertz frequency to optical frequency, offer an appealing solution for low-noise photonic microwave and millimeter-wave generation Liu et al. (2020); Zhang et al. (2019); Wang et al. (2021); Tetsumoto et al. (2021); Yao et al. (2022). Photodetection of the microcomb’s pulse stream creates a microwave whose carrier frequency corresponds to the microcomb’s repetition rate $f_{\text{rep}}$. Several methods have been implemented to further improve the microwave’s spectral purity, including the use of an external microwave Weng et al. (2019), an auxiliary laser Liu et al. (2024), a transfer comb Lucas et al. (2020), or two-point optical frequency division Kudelin et al. (2024); Sun et al. (2024b); Zhao et al. (2024).
@@
Here we harness the platicon switching dynamics, and demonstrate a unique noise-quenching effect allowing suppression of microwave’s phase noise. In contrast to methods mentioned above, our system is free-running, all passive (i.e. without any active locking), and elegantly simple (i.e.without extra RF or laser sources). Experimentally, we collect the output light from the microresonator with a commercial photodetector (PD). The PD converts the dark pulse stream of $f_{\text{rep}}=10.7$ GHz to a 10.7-GHz microwave in the critical X-band which is dedicated for radar, wireless networks, and satellite communication.
@@
By varying the laser current $I$ within the SIL regime, we monitor the platicon’s $f_{\text{rep}}$ and phase noise $S_{\phi}$ at 10 kHz Fourier offset frequency of the 10.7-GHz microwave, as shown in Fig. 3d. For each step, a local minimum of $S_{\phi}$ is observed. Meanwhile, it coincides with the local maximum of $f_{\text{rep}}$, as highlighted with vertical dashed lines in Fig. 3d. This coincidence is due to that, at the local maximum of $f_{\text{rep}}$, we have $\text{d}f_{\text{rep}}/\text{d}I=0$, as highlighted with horizontal dashed lines in Fig. 3d. Therefore, $f_{\text{rep}}$ becomes insensitive to current-noise-induced laser frequency jitter, resulting in the lowest $S_{\phi}$. More analysis of this coincidence is found in Supplementary Information Note 4. In fact, this effect is similar to the “quiet point” effect observed in bright dissipative solitons without SIL Yi et al. (2017); Triscari et al. (2023). The underlying mechanism is attributed to the multi-mode coupling between the laser cavity and the Si3N4 microresonator, which causes asymmetric comb-line enhancement or suppression, as shown in Fig. 3a. As a result, $f_{\text{rep}}$ depends on the laser frequency (be more specifically, the detuning of the free-running laser frequency to the cold resonance frequency). In our theoretical model, only the pump resonance is considered coupled, i.e. $b_{0}$ in Eq. (2). A more precise simulation on the optical spectra should include multi-mode coupling by substituting $b_{0}$ with $\sum_{\mu}b_{\mu}e^{i\mu d_{1}(\tau-\tau_{\text{s}})}$.
@@
The $S_{\phi}$ spectra of the local maximum and minimum points, marked with green and black dots in Fig. 3d, are measured and compared in Fig. 3e, showing 23.5 dB noise reduction. The lowest $S_{\phi}$ reaches 45/75/105-45/-75/-105- 45 / - 75 / - 105 dBc/Hz at 0.1/1/10 kHz Fourier offset frequency, and is limited by the shot noise floor at higher frequency offset Savchenkov et al. (2008). Figure 3e inset shows the beatnote of the lowest $S_{\phi}$ (black data) with 10 Hz resolution bandwidth (RBW). Supplementary Information Note 5 illustrates that, such a noise quenching phenomenon has been also observed in parallel experiments with different Si3N4 microresonators and DFB lasers, thus is universal.
@@
In conclusion, we unveil an intriguing yet universal Kerr-thermal dynamics of a semiconductor laser self-injection-locked to a Si3N4 microresonator where platicon microcombs are formed. We experimentally observe and characterize the platicon switching dynamics with discrete steps. We further establish a comprehensive theoretical model cooperating the laser-cavity dynamics, the platicon formation dynamics, the photo-thermal dynamics, and the mutual coupling between them. Numerical simulation confirms the experimental result, and illuminates that the platicon switching phenomenon is originated from the synergy of SIL, Kerr nonlinearity and the photo-thermal effect. Exploiting this finding, we showcase low-noise microcomb-based microwave generation. Via operation of platicons with specific laser current, we achieve 23.5 dB phase noise quenching of the 10.7-GHz microwave carrier. Our study not only add critical insight of pulse formation in linear-and-nonlinear-coupled laser-microresonator systems, but also offer a neat solution for photonic-chip-based microwave oscillators with high spectral purity, ideal for microwave photonics, coherent optical communication, analog-to-digital conversion, wireless links, and radar.
@@
Acknowledgments: We thank Baoqi Shi for characterizing the Si3N4 chips, and Zhiyang Chen for assistance in the experiment. We acknowledge support from the National Natural Science Foundation of China (Grant No.12261131503, 12404436), Innovation Program for Quantum Science and Technology (2023ZD0301500), Guangdong-Hong Kong Technology Cooperation Funding Scheme (Grant No. 2024A0505040008), Shenzhen-Hong Kong Cooperation Zone for Technology and Innovation (HZQB-KCZYB2020050), and Shenzhen Science and Technology Program (Grant No. RCJC20231211090042078). Russian Quantum Center was supported by RSF grant (23-42-00111). The DFB lasers were fabricated by Shenzhen PhotonX Technology Co. Ltd. and Henan Shijia Photons Technology Co. Ltd. . Silicon nitride chips were fabricated by Qaleido Photonics.

### v2

@@
We further perform numerical simulation of Eqs. (1–5) using experimentally obtained or realistic values for each parameter. Details are found in Supplementary Information Note 3. Figure 2(d,e) presents the simulated output optical power and frequency, which agree with the experimental data on both the overall trend (due to the photo-thermal effect) and the platicon steps (due to the Kerr nonlinearity). Figure 2f shows the zoom-in profile of the gray-shaded zone in Fig. 2(d,e), and highlights the full detuning range of platicon steps with gray curves. Steps corresponding to those in Fig. 2c are marked. We emphasize that, previous efforts [41, 33, 24] to model platicon formation in the SIL regime fail to investigate and analyze the transient behaviour, and thus fail to reveal the step features. Meanwhile, they have also omitted the photo-thermal effect. Although the Lugiato-Lever equation (LLE) can capture platicon steps,it does not contain the SIL dynamics. Detailed comparison between our model and the LLE model is found in Supplementary Information Note 4.
@@
In Fig.3a, the red curves presents the simulated platicon spectra for each labeled steps in Fig. 2f. They confirm the experimental data, not only on the spectral envelopes but also on the number of fringes (marked with arrows). Figure 3a insets present the simulated time-domain pulse shapes, revealing that the fringes are related to the oscillating tails (marked with arrows). Unlike the steps of bright dissipative solitons [42, 31], we demonstrate here that platicon steps can occur even in the single-platicon state, i.e. only one dark pulse in the microresonator.
@@
Low-noise photonic microwave generation. Finally, we showcase an immediate application of our findings, i.e. to generate low-noise photonic microwave. Photonic microwave – microwave synthesized via photonics – allows unrivalled noise performance and bandwidth breaking the bottleneck of their electronic counterparts [43, 44, 45]. Microcombs offer an appealing solution for low-noise microwave and terahertz generation [46, 47, 48, 49, 50]. Photodetection of the microcomb’s pulse stream creates a microwave whose carrier frequency corresponds to the microcomb’s repetition rate $f_{\mathrm{rep}}$. Several methods have been implemented to further improve the microwave’s spectral purity, including the use of an external microwave [51], an auxiliary laser [52], a transfer comb [53], or two-point optical frequency division [54, 55, 56].
@@
Here we harness the platicon switching dynamics, and demonstrate a unique effect allowing suppression of the microwave’s phase noise. In contrast to the methods mentioned above, our system is free-running, all passive (i.e. without any active locking), and simple. Experimentally, we collect the output light from the microresonator with a commercial photodetector (PD). The PD converts the dark pulse stream of $f_{\mathrm{rep}}=10.7$ GHz to a 10.7-GHz microwave in the critical X-band.
@@
By varying the laser current $I$ within the SIL regime, we monitor the platicon’s $f_{\mathrm{rep}}$ and phase noise $S_{\phi}$ at 10 kHz Fourier offset frequency of the 10.7-GHz microwave, as shown in Fig. 3b. For each step, a local minimum of $S_{\phi}$ is observed. Meanwhile, it coincides with the local maximum of $f_{\mathrm{rep}}$, as highlighted with vertical dashed lines in Fig. 3b. This coincidence is due to that, at the local maximum of $f_{\mathrm{rep}}$, we have $\mathrm{d}f_{\mathrm{rep}}/\mathrm{d}I=0$, as highlighted with horizontal dashed lines in Fig. 3b. Therefore, $f_{\mathrm{rep}}$ becomes insensitive to current-noise-induced laser frequency jitter, resulting in the lowest $S_{\phi}$. More analysis of this coincidence is found in Supplementary Information Note 5. In fact, this effect is similar to the “quiet point” effect observed in bright dissipative solitons without SIL [57, 58]. The underlying mechanism is attributed to the multi-mode coupling between the laser cavity and the Si3N4 microresonator, which causes asymmetric comb-line enhancement or suppression, as shown in Fig. 3a. As a result, $f_{\mathrm{rep}}$ depends on the laser frequency (be more specifically, the detuning of the free-running laser frequency to the cold resonance frequency). In our theoretical model, only the pump resonance is considered coupled, i.e. $b_{0}$ in Eq. (2). A more precise simulation on the optical spectra should include multi-mode coupling by substituting $b_{0}$ with $\sum_{\mu}b_{\mu}\mathrm{e}^{i\mu d_{1}(\tau-\tau_{\mathrm{s}})}$.
@@
The $S_{\phi}$ spectra of the local maximum and minimum points, marked with green and black dots in Fig. 3b, are measured and compared in Fig. 3c, showing 23.5 dB noise reduction. The lowest $S_{\phi}$ reaches 45/75/105-45/-75/-105- 45 / - 75 / - 105 dBc/Hz at 0.1/1/10 kHz Fourier offset frequency, and is limited by the shot noise floor at higher offset frequency[59]. Figure 3c inset shows the beat note of the lowest $S_{\phi}$ (black data) with 10 Hz resolution bandwidth (RBW). Supplementary Information Note 6 illustrates that, the identical phenomenon has been also observed in parallel experiments with different Si3N4 microresonators and DFB lasers, and thus is universal.
@@
In conclusion, we unveil an intriguing yet universal Kerr-thermal dynamics of a semiconductor laser self-injection-locked to a Si3N4 microresonator where platicon microcombs are formed. We experimentally observe and characterize the platicon switching dynamics with discrete steps. We further establish a comprehensive theoretical model cooperating the laser-cavity dynamics, the platicon formation dynamics, the photo-thermal dynamics, and the mutual coupling between them. Numerical simulation confirms the experimental result, and illuminates that the intriguing rich platicon switching dynamics is originated from the synergy of SIL, Kerr nonlinearity and the photo-thermal effect. Exploiting this finding, we showcase low-noise platicon-based microwave generation, and achieve 23.5 dB phase noise reduction of the 10.7-GHz microwave carrier. Our study not only adds critical insight into pulse formation in linear-and-nonlinear-coupled laser-microresonator systems, but also offers a neat solution for photonic-chip-based microwave oscillators with high spectral purity, ideal for microwave photonics, coherent optical communication, analog-to-digital conversion, wireless links, and radar.
@@
Acknowledgments: We thank Baoqi Shi for characterizing the Si3N4 chips. We acknowledge support from the National Natural Science Foundation of China (Grant No.12261131503, 12404436), Innovation Program for Quantum Science and Technology (2023ZD0301500), National Key R&D Program of China (Grant No. 2024YFA1409300), Guangdong-Hong Kong Technology Cooperation Funding Scheme (Grant No. 2024A0505040008), Shenzhen-Hong Kong Cooperation Zone for Technology and Innovation (HZQB-KCZYB2020050), and Shenzhen Science and Technology Program (Grant No. RCJC20231211090042078). Russian Quantum Center was supported by RSF grant (23-42-00111). The DFB lasers were fabricated by Shenzhen PhotonX Technology Co. Ltd. and Henan Shijia Photons Technology Co. Ltd. . Silicon nitride chips were fabricated by Qaleido Photonics.